A person has a mass of 60.0 kg. She is standing on a scale in an elevator that is at rest. The elevator starts accelerating upward at a rate of 2.10 m/sec2.
(a)How far has the elevator moved upward at t = 3.15 seconds?
(b)What is the speed of the elevator at t = 3.15 seconds?
1
Expert's answer
2020-11-24T06:19:44-0500

Given,

Mass of person="60kg"

Accelaration of elevator "a=2.10m\/s^2"

Initialy It is at rest so "u=0m\/s"


a) at "t=3.15s"


Apply second equation of motion,

"s=ut+\dfrac{1}{2}at^2
\\\Rightarrow s=\dfrac{1}{2}(2.10)(3.15)^2\\\Rightarrow s=10.5(9.925)=104.18m"


Hence Distance covered by lift is"104.18m"


(b)Apply first equation of motion

Let the speed at "3.15s" be "v"


"v=u+at
\\\Rightarrow v=0+(2.10)(3.15)=6.615m/s"


Hence the velocity is "6.615m\/s"